Circular No. 7833

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE A. Becker, Bell Laboratories, Lucent Technologies, reports, on behalf

of the Deep Lens Survey Team (cf. IAUC 7804) and P. Baca (University of Colorado at Denver), the discovery of eight apparent supernovae in images taken on the Cerro Tololo 4-m telescope (+ MOSAIC II imager) by I. Dell’Antonio, Baca, D. Norman, and Becker.

SN 2002 UT α_{2000} δ_{2000} Mag.
2002ax Feb. 14.3 10^{h}48^{m}24.^{s}43 − 4^{o}11′27′′.7 21.2 V
2002ay Feb. 14.3 10 49 28.01 − 4 29 43.6 21.6 V
2002az Feb. 14.2 10 54 17.44 − 4 54 16.0 22.9 z′
2002ba Feb. 15.0  5 14 37.33 −49 07 44.4 23.6 V
2002bb Feb. 15.0  5 15 06.05 −48 59 51.1 24.1 B
2002bc Feb. 15.0  5 16 16.63 −48 53 23.0 23.8 R
2002bd Feb. 15.3 10 55 35.71 − 5 42 04.6 23.1 R
2002be Feb. 16.3 10 49 54.81 − 5 46 12.5 24.1 R

The new objects were identified by differencing five stacked exposures from previous sets of five stacked exposures (whose limiting magnitudes are listed below). Additional magnitudes: SN 2002ax, 2002 Jan. 11.3, V ∼ [25.5; Feb. 16.3, B = 21.9. SN 2002ay, 2002 Jan. 11.3, V ∼ [25.5; Feb. 16.3, B = 23.3. SN 2002az, 2002 Mar. 4.4, z′ ∼ [25.5; 2002 Feb. 16.2, B = 23.7; 17.2, R = 21.2. SN 2002ba, 2002 Jan. 10.1, R ∼ [26; Feb. 19.1, V = 22.3. SN 2002bb, 2002 Jan. 10.1, R ∼ [26; Feb. 19.1, V = 21.1. SN 2002bc, 2002 Jan. 10.1, R ∼ [26; Feb. 19.1, V = 21.7. SN 2002bd, 2002 Jan. 10.2, V ∼ [25.5; Feb. 18.2, R = 22.26. SN 2002be, 2002 Jan. 15.2, B ∼ [25.5; Feb. 17.2, R = 23.6; 19.3, V = 23.4.

SUPERNOVA 2001jx IN NGC 4622 R. Buta and G. G. Byrd, University of Alabama; and T. Freeman,

Bevill State Community College, report their discovery of an apparent supernova (U = 17.5) on Hubble Space Telescope images taken with filters F336W (U), F439W (B), F555W (V), and F814W (I) during 2001 May 25.6–25.9 UT. SN 2001jx is located at α = 12^{h}42^{m}40^{s}.04, δ = −40^{o}43′12′′.5 (equinox 2000.0), which is 28′′.6 east and 87′′.4 north of the nucleus of NGC 4622. There is nothing at this position on the Digitized Sky Survey. During the 8-hr observing period, there was no detectable movement of the object. At the time of the observations, the object was bright enough to saturate the B, V, and I images in 500-s exposures.

2002 February 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green